OpenStudy (anonymous):

if the value of asintheta +bcostheta=c then the value of acostheta-bcostheta is

OpenStudy (cp9454):

first square both the sides then convert sintheta to costheta and solve

OpenStudy (anonymous):

\[(a \sin \theta)^2+(b \cos \theta)^2=c^2\]

OpenStudy (anonymous):

(asinx+bcosx)^2=c^2

OpenStudy (cp9454):

no squaring is not done like that you have to do the whole square

OpenStudy (cp9454):

\[(asin \theta+bcos \theta)^{2}\]

OpenStudy (anonymous):

a^2sin^2x+b^2cos^2x+2asinxbcox=c^2

OpenStudy (anonymous):

for convenience i am using x

OpenStudy (cp9454):

yup like that, and then change sintheta into costheta and solve

OpenStudy (anonymous):

a-acos^2x+bcos^2x+2absinxcosx=c^2 a-cos^2x(a+b)+2absinxcosx=c^2

OpenStudy (anonymous):

sorry in my question acostheta-bsinthetha to find

OpenStudy (cp9454):

i am telling u wait.

OpenStudy (cp9454):

i have solved in the attachment

OpenStudy (anonymous):

i have many doubts

OpenStudy (cp9454):

you have seen the attachment